Periodic Inventory by Three Methods; Cost of Merchandise Sold The units of an item available for sale during the year were as follows:

Jan. 1 Inventory 40 units @ $108

Mar. 10 Purchase 70 units @ $116

Aug. 30 Purchase 20 units @ $120

Dec. 12 Purchase 70 units @ $126

There are 80 units of the item in the physical inventory at December 31. The periodic inventory system is used.

Determine the inventory cost and the cost of merchandise sold by three methods. Round interim calculations to one decimal and final answers to the nearest whole dollar. Cost of Merchandise Inventory and Cost of Merchandise Sold

Inventory Method: Merchandise Inventory: Merchandise Sold

First-in, first-out (FIFO):

Last-in, first-out (LIFO):

Weighted average cost:
Answer #1
Units Unit cost Total
Jan 1 40 108 4320
Mar 10 70 116 8120
Aug 30 20 120 2400
Dec 12 70 126 8820
Total 200 23660
Average cost per unit 118.3 =23660/200
First-in, first-out (FIFO):
Merchandise Inventory 10020 =(70*126)+(80-70)*120
Merchandise Sold 13640 =23660-10020
Last-in, first-out (LIFO):
Merchandise Inventory 8960 =(40*108)+(80-40)*116
Merchandise Sold 14700 =23660-8960
Weighted average cost:
Merchandise Inventory 9464 =80*118.3
Merchandise Sold 14196 =23660-9464